Let us assume that the probability of winning a lottery ticket is 5x10-6. Find the expected number of tickets you need to buy to win the lottery.
Answer #1

here as we know that expected number of tickets you need to buy to win the lottery =1/p

=1/(5*10-6)=0.2*106 =200000